OpenStudy (anonymous):

The equation of a line is. What will the slope be for a line that is perpendicular to this line? m=-2/7 m= 2/7 m=7/2 m=-7/2

hartnn (hartnn):

equation of line is?

OpenStudy (anonymous):

To make a slope perpendicular, you just take the reciprocal on the opposite. In other words, you flip the fraction, and multiply by -1. so -2/1 becomes 1/2

OpenStudy (anonymous):

the line is y=2/7x-5. sorry i forgot to put that in there!!!

hartnn (hartnn):

compare it with y=mx+c and tell me what is m(slope)

OpenStudy (anonymous):

the slope is 2/7

hartnn (hartnn):

yes

hartnn (hartnn):

now product of slope of perpendicular lines is -1 so m(2/7) =-1 can u find m from here

OpenStudy (anonymous):

so it would b -1x-5?

hartnn (hartnn):

what?

hartnn (hartnn):

m(2/7) =-1 m= -7/2

OpenStudy (anonymous):

ooh thanks ;D

hartnn (hartnn):

welcome ^_^